OpenStudy (anonymous):

how do you solve this differential equation? (y+1)dx+(4x-y)dy=0 ??

OpenStudy (anonymous):

I was going to say split it up (y+1)dx+(4x-y)dy=0 dx/(4x-y)=-dy/(y+1)

OpenStudy (anonymous):

But you are still left with a y in the left side.

OpenStudy (anonymous):

yeah xD

OpenStudy (anonymous):

Sorry just throwing out ideas hoping you might catch something.

OpenStudy (lgbasallote):

@shinigami1m you can use exact differential equation...do you know how to use that?

OpenStudy (lgbasallote):

wait i mean integrating factor for exact differential equation

OpenStudy (anonymous):

like partial M and partial N?

OpenStudy (lgbasallote):

you can also use the bloody and morbid homogeneous

OpenStudy (anonymous):

Oh is this when you take a derivative of each one plug it in to each other integrate and so on?

OpenStudy (lgbasallote):

@shinigami1m this is what im talking about...it's not discussed in most schools so i dont know if you know it... http://openstudy.com/updates/4fd16ab5e4b057e7d220ae39

OpenStudy (anonymous):

Gj on the tutorial @lgbasallote

OpenStudy (anonymous):

i know that method :D

OpenStudy (anonymous):

i got (y+1)^3 for integrating factor

OpenStudy (lgbasallote):

wait..why? \[\frac{\partial M}{\partial y} - \frac{\partial N}{\partial x} = 4 - 1 = 3\] right?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

wait i mean -3

OpenStudy (anonymous):

umm its 1-4

OpenStudy (lgbasallote):

it's 1- 4

OpenStudy (lgbasallote):

then you divide it by M so you have \[-\frac{3}{y+1}\] right? btw thanks @Romero

OpenStudy (anonymous):

yep

OpenStudy (lgbasallote):

then you use \[e^{-\int g(y)dy}\] right?

OpenStudy (anonymous):

yeeeee

OpenStudy (lgbasallote):

so you have \[\huge e^{\int \frac{3}{y+1}dy}\] agree?

OpenStudy (anonymous):

agree

OpenStudy (lgbasallote):

lol yeah it's (y+1)^3 sorry :p haha

OpenStudy (anonymous):

HAHA the hard part is the next part... :| distributing I.F.!!

OpenStudy (lgbasallote):

so i guess using homogeneous is easier in this one :/

OpenStudy (anonymous):

its not homogenous i think

OpenStudy (lgbasallote):

hmm yeah...darn y+1 -_-

OpenStudy (lgbasallote):

how about linear DE?

OpenStudy (lgbasallote):

it's linear in y

OpenStudy (lgbasallote):

\[\frac{dx}{dy} + (\frac{4}{y+1})x = \frac{y}{y+1}\]

OpenStudy (anonymous):

i'm at the final part... i just need to integrate -y(y+1)^3dy~~~~

OpenStudy (lgbasallote):

lol this is nice...there are a lot of ways to solve this... which method did you use?

OpenStudy (anonymous):

exactness i just continued it haha... i think linearity is easier

OpenStudy (lgbasallote):

yeah it is lol

OpenStudy (lgbasallote):

IF = (y+1)^4 i think

OpenStudy (anonymous):

that is if we use linearity haha y(y+1)^3 on the other side still XD but i got the answer hhehe

OpenStudy (anonymous):

I always like youtubing these to see how other people do it. I like this one http://www.youtube.com/watch?v=bwASJWS8ltM You know I hear that you are able to retain information better if you do more study more than one way. Like reading and hearing other people then you doing it by yourself.

OpenStudy (anonymous):

So I hope that helps.